OpenStudy (anonymous):

A team wins 105 games. This is 70% of the games played. How many games were played?

OpenStudy (anonymous):

150

OpenStudy (saifoo.khan):

how?

OpenStudy (anonymous):

105 = 70% * x 105 = 0.70 * x x = 105/0.70 x = 150

OpenStudy (gw2011):

Let x = games played 0.70x = 105 x = 105/0.70 x = 150